OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -2) and (-1,0)? -1/2 1/2 2 -2 I end up getting 2/-4 in that right

OpenStudy (anonymous):

Slope is defined as \[m = {rise \over run} = {y_1 - y_0 \over x_1 - x_0}\]By this definition, the answer you got is correct.

OpenStudy (anonymous):

m = [ -2- 0 ] / [ 3- (-1) ] = -2 / 4 = - 1/2